# Collimated $\gamma$-flash emission along the target surface irradiated by a laser at non-grazing incidence

### v1

@@ Abstract
The interaction of a high-power laser with a solid target provides ways to produce beams of $\gamma$-photons. For normal incidence of the laser on the target the beams usually appear in a form of two lobes, which are symmetric with respect to the laser propagation axis. In this work we demonstrate via three-dimensional particle-in-cell simulations a regime where for oblique incidence the emission of a collimated $\gamma$-photon beam is in the direction parallel to the target surface. The process is ascribed to the interference pattern in the electromagnetic field formed by the incident and reflected laser pulse. The electromagnetic field accelerates electrons to the GeV energy level, while temporarily directing their momentum along the target surface. Consequently, they emit a collimated $\gamma$-photon beam in the same direction. The dependencies of $\gamma$-photon emission on the incident angle, laser pulse polarization, power and duration and target thickness are also addressed in the paper. The beam directionality is important for designing future experiments. In addition, this setup causes the generation of high-order harmonics propagating along the target surface.
  • October 16, 2024
@@ 1 Introduction
An enormous progress in the development of extremely high-power laser technology resulting in the construction of many multi-PW laser facilities has been made all over the world [1]. As a result, there is a fast growth of interest towards building compact laser-based accelerators of charged particles [2, 3, 4, 5, 6, 7, 8] and radiation sources [9, 10]. Recent experiments on laser electron acceleration demonstrated electrons accelerated up to the energy of 10 gigaelectron volts (GeV) within a distance of tens of centimeters [11, 12]. In ion acceleration, proton beams with energy reaching 150 MeV have been demonstrated in the interaction of a laser with a thin foil target [13]. High-power lasers are used for studying high-brightness high-energy photon sources [14] and high-order harmonic generation (HHG) in the laser interaction with gas [15] and solid density targets [9, 16]. These lasers pave the way towards studying the previously unexplored domain of parameters corresponding to superstrong field for the quantum regimes of electromagnetic wave interaction with charged particles and the vacuum [17, 18, 19, 20]. These processes have been demonstrated through the electron-positron ($e^{-}$$e^{+}$) pair creation from vacuum in the seminal experiment [21]. A further experiment has been recently proposed for proving the nonlinear property of vacuum [22]. It was also predicted theoretically that through this interaction the $\gamma$-flash can be generated via multiphoton Compton scattering with high conversion efficiency [23, 24, 25, 26] and the first experimental evidence for this has been recently presented [27]. The multiphoton Compton scattering is characterized by the Lorentz invariant parameter $\chi_{e}=\gamma_{e}\sqrt{\left(\textbf{E}+\textbf{v}\times\textbf{B}\right)^{2}-\left({\textbf{v}\cdot\textbf{E}/c}\right)^{2}}$, where $\gamma_{e}$ is the Lorentz factor of the electron with velocity v, $c$ is the speed of light in vacuum, and E and B are the laser electric and magnetic fields normalized to the Schwinger field, $E_{S}\approx 1.3\times 10^{18}$ $\mathrm{V/m}$. The $\gamma$-flash radiation has exciting applications in various fields, such as materials science at extreme energy densities [28], photonuclear reactions [29, 30, 31], neutron sources [32], photonuclear fission [33, 34], radiotherapy [35], shock-wave studies [36], electron-positron ($e^{-}$$e^{+}$) pair generation [37] (which can be further guided using orthogonal collision with a laser [38, 39]) and in quantum technologies[40], and it can be used to help understand the mechanisms of high-energy astrophysical processes [41, 42, 43, 44].
@@
A typical approach towards realization of $\gamma$-photon radiation uses the interaction of a laser pulse with a solid density target [45, 46, 47, 26]. Here, the laser pulse accelerates electrons, which then interact with the same pulse. Another setup to increase the $\gamma$-photon emission is to use two laser pulses [48, 49, 50, 51, 52] (or more of them [53, 54]). The normal incidence of a linearly polarized laser on target results in a symmetric double-lobe structure of the emitted $\gamma$-photons, which is demonstrated in particle-in-cell (PIC) simulations for various cases [24, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. However, in typical laser-solid interaction experiments, an oblique incidence angle is required to mitigate laser back-reflection which could damage the laser. In contrast to the double-lobe of the normal incidence, the structure and directionality of the emitted $\gamma$-photon beam(s) for oblique incidence varies strongly for different configurations. In addition to the incidence angle, the $\gamma$-photon directionality also depends on the target density, surface modulation [65] and target thickness, and various laser parameters (e.g., duration and focal spot size) [66]. Thus, the prediction of $\gamma$-photon beam direction becomes crucial for the reliable measurement of future experiments. A possible solution is to use specific laser-target parameters to align the dominant direction of the emitted $\gamma$-photons to the target surface. This was previously reported in simulations using grazing laser incidence in combination with a relativistically near-critical target [67] (although limiting the $\gamma$-photon yield) and in simulations focusing on lower-energy radiations (extreme ultraviolet - XUV for oblique incidence [16] and hard x-ray for parallel incidence [68]). For oblique incidence, interference between incident and reflected parts of the laser pulse generates a pattern in the electromagnetic field, accelerating electrons to high energies, as demonstrated via grazing incidence [69, 70]. This approach can be extended to a non-grazing incidence and opaque target (where the generated $\gamma$-flash is separated from the reflected laser pulse) as demonstrated in this work. The electron momentum is temporarily directed along the target surface by the field pattern, which results in a collimated $\gamma$-photon beam being emitted in the same direction.
@@
The paper is structured as follows. The theory of electron dynamics in the interference pattern of the electromagnetic field is described in Section 2. The simulation method and parameters are described in Section 3. The results are presented in Section 4, which is divided into four subsections. In subsection 4.1 we compare the normal incidence with ${}^{\circ}$oblique incidence and describe the mechanism behind the collimated $\gamma$-photon emission in the direction parallel to the target surface. This section also includes a Fourier transform analysis to observe the HHG. In the following subsections the dependence of the $\gamma$-photon emission on different laser and target parameters is discussed. Specifically, the dependence on the laser incidence angle is shown in subsection 4.2, on the laser polarization in subsection 4.3 and finally the dependence on the laser duration, laser power and target thickness in subsection 4.4. Section 5 contains the discussion of the results and the conclusion.
@@ 3 Simulation setup and parameters
The 3D simulations were done with the PIC code EPOCH [71], including the Higuera-Cary pusher and QED module for non-linear Compton scattering. The Bremsstrahlung process [72, 73] is omitted in this work, as Compton scattering dominates over it at ultra-high laser intensities for times comparable to the laser pulse duration. The simulation box size is 14.336 $\mu$m $\times$ 20.48 $\mu$m $\times$ 10.24 $\mu$m, the cell size is 4 nm $\times$ 4 nm $\times$ 32 nm and 4 macro-electrons and 4 macro-ions are assigned to each cell containing plasma. In all simulations the laser propagates from the boundaries at positions $x=y=-10.24$ $\mu$m at an angle of 45 (the plane of incidence is the $x$-$y$ plane), as shown in figure LABEL:fig:sim_setup. The incidence angle is determined by the target orientation, as shown in Fig. LABEL:fig:sim_setup(a) for the normal incidence and LABEL:fig:sim_setup(b) for the 45 incidence.
@@
The reference laser is linearly p-polarized. A comparison with linear s-polarization and circular polarization is discussed in Section 4.3. The laser wavelength is $0.8\ \mathrm{\mu m}$ and the peak intensity is $I_{\mathrm{max}}=2\times 10^{23}\ \mathrm{W/cm^{2}}$, thus $a_{0}\approx 0.85\sqrt{I\left[10^{18}\mathrm{W\ cm^{2}}\right]\lambda^{2}\left[\mathrm{\mu m}\right]}\approx 304$. The critical plasma density is $n_{c}=~{}m_{e}\omega^{2}/4\pi e^{2}\approx 1.742\times 10^{21}\ \mathrm{cm^{-3}}$, where $\omega$ is the laser angular frequency, $m_{e}$ is the electron mass and $e$ is the electron charge. The laser pulse has Gaussian spatial and temporal profiles with beam width equal to 3 $\lambda$ (2.4 $\mu$m) at the full width at half maximum (FWHM) and duration of 20 fs at FWHM (in intensity). The total laser energy for the reference case is 283 J in the 3D simulation, thus the laser power is approximately 14 PW. Aside of the default parameters, Section 4.4 compares lasers of different intensities (ranging from $2\times 10^{22}\ \mathrm{W/cm^{2}}$ to $10^{24}\ \mathrm{W/cm^{2}}$ ) and pulse durations (ranging from 15 to 30 fs). The target is made of fully ionized iron with an electron number density of $2.22\times 10^{24}\ \mathrm{cm}^{-3}$ (about 1274 $n_{c}$). The target thickness is 1.5 $\mu$m (a target thickness comparison is made in Section 4.4).
Photons are set to be immobile, which prevents them from escaping the simulation box and also reduces the computational time. Furthermore, only $\gamma$-photons with energy above 1 MeV are stored (photons with lower energy still contribute to the electron recoil). The exceptions are the two simulations shown in Fig. LABEL:fig:sim_setup, where photon movement is enabled in order to visualize their dynamics. These simulations are also used to test the significance of $e^{-}$$e^{+}$ pair production for our simulation parameters. However, only one positron macroparticle is generated for the ${}^{\circ}$case and 4 for the ${}^{\circ}$case.
The visualization of the $\gamma$-photon emission temporal evolution of the ${}^{\circ}$case can be accessed online [74] via our Virtual Beamline – VBL application [75, 76, 77]. VBL uses a custom-made WebGL [78] application to render the visualization inside a regular web browser, as shown in Fig 2. The application also enables support for Virtual Reality (VR) viewing, which was tested with Oculus Rift S and HP Reverb G2 headsets.
Figure 2: Screenshot of the interactive VBL visualization of the ${}^{\circ}$incidence case.
@@ 4.1 Generation of collimated $\gamma$-photon beam along the target surface
The $\gamma$-photon directionality is shown in Fig. LABEL:fig:spherical_distribution_00x45 via the angular distribution of the emitted $\gamma$-photon energy in the 3D space. Due to the setup geometry with the $x$-$y$ polarization plane (whose 2D slices at $z=0$ are further examined later), it is convenient to represent the angular distribution using the angles $\theta=\mathrm{arctan}(p_{y}/p_{x})$ and $\phi=\mathrm{arctan}(p_{z}/p_{r})$ in Fig. LABEL:fig:spherical_distribution_00x45, where $p_{x,y,z}$ are the momenta in the respective directions of the Cartesian system of the simulation and $p_{r}=\sqrt{p_{x}^{2}+p_{y}^{2}}$.
Note that as targets are tilted by a different angle for different laser incidence angles, the angular data are shifted in the $\theta$-direction by an angle equal to the angle between the target and the $y$-axis of the simulation. The normal incidence case (Fig. LABEL:fig:spherical_distribution_00x45-a) creates a typical double-lobe structure. For our simulation parameters the maximum energy density is observed around ±$\pm 50^{\circ}$. The situation differs for the 45 case (figure LABEL:fig:spherical_distribution_00x45-b) where a single spot around 90 (direction parallel to the target surface) is observed. The lineouts of the blue lines (right part of Fig. LABEL:fig:spherical_distribution_00x45) show the energy distribution in the transverse direction at angles $50^{\circ}$ and $90^{\circ}$ for the normal and ${}^{\circ}$incidence cases respectively. In both cases the $\gamma$-photon beam is well collimated in the transverse direction (angular spread at FWHM is around ${}^{\circ}$for both cases) but the maximal energy density is more than double in the case of oblique incidence. The brilliance in Fig. LABEL:fig:spherical_distribution_00x45-c is calculated at the direction of the peak $\gamma$-photon yield (at the same angles as the lineouts) using an opening full-angle of 10. Our simulations reveal a high brilliance for the ${}^{\circ}$case, similar to that of Ref. [62], approaching $10^{23}$ s-1mm-2mrad-2 per 0.1% bandwidth in the tens of MeV $\gamma$-photons energy range. In comparison, the brilliance for the normal incidence case is approximately three times lower, which is reflected to the ratio of the conversion efficiency, as discussed in Section 4.2.
@@
For the $45^{\circ}$ laser incidence case (shown in Fig. LABEL:fig:field_electron_gamma-b,d) $\gamma$-photons are emitted by three distinct mechanisms, occurring at different locations and with different electron trajectory patterns. Firstly, as described in Section 2, the reflected laser pulse interferes with the incident one and creates an interference pattern, which propagates with superluminal phase velocity along the target surface (as seen in Fig. LABEL:fig:field_electron_gamma-b). Inside this pattern, electrons are accelerated to energies several times higher than that given by the quiver velocity [70]. As the field pattern moves faster than electrons, they experience a strong acceleration (and subsequent deceleration when the pattern changes sign) while temporarily aligning their momentum with the target surface. The direction of motion of the emitted $\gamma$-photons is approximately the same as that of the emitting electrons. Therefore, directional $\gamma$-photon emission occurs at the field extremes, as in Figs. LABEL:fig:field_electron_gamma-b,d. The field extremes and the regions of high $\gamma$-photon energy density appear at $y=[-1.7,-0.6,0.55,1.7]$ $\mu\mathrm{m}$, with their spacing as predicted by the equation $\lambda_{\perp}=\lambda/\sin\theta_{n}=0.8\ \mu\mathrm{m}/\sin(45^{\circ})\approx 1.13\ \mu\mathrm{m}$. These spots appears around $x=0$, which at time 60 fs is a low-density plasma region in front of the target (as the target surface shifts inwards under the radiation pressure). The instantaneous target surface is shown by the green contour in Fig. LABEL:fig:field_electron_gamma-b.
@@
Thirdly, with an even smaller contribution, a $\gamma$-photon population is emitted from the instantaneous target surface ($y=[-1,0.25]$ $\mu\mathrm{m}$ at 60 fs). There, electrons propagate along the target surface with wiggling movement, as seen in Fig. LABEL:fig:field_electron_gamma-b.
@@
Here, $C=e^{2}m_{e}^{2}/3\sqrt{3}\pi\hbar^{2}$, where $\hbar$ is normalized Planck constant, $K_{2/3}$ is the Modified Bessel function of the second kind and $u$ is the the integration variable. Note that Fig. LABEL:fig:trajectory-a is similar to Fig. LABEL:fig:field_electron_gamma-d (although Fig. LABEL:fig:trajectory-a is more clear since the electron data are instantaneous).
@@
Next, we tracked electrons with $\chi_{e}\geq 0.3$ (34 macroparticles), located in the bright spot area shown in Fig. LABEL:fig:trajectory-a. The electron trajectories are shown in Fig. LABEL:fig:trajectory-b, where their energy is represented by the color scale (time evolution of the location of a selected particle in $B_{z}$, $E_{x}$, $E_{y}$ is shown in the supplementary video). The evolving (with respect to the $y$-position) front target surface is depicted by the green line, the $x$-position represents the first position with electron number density equal or higher than the initial target density at $y=z=0$ and the $y$-position is linked to one of the tracked electrons. The tracked electrons begin with a wiggling movement close to the front target surface as they reflect from the strong quasi-static magnetic field (generated by the space charge inhomogeneity). Afterwards, they are injected into the field interference pattern, align parallel to the initial target surface, and are accelerated to energies up to 700 MeV. Consequently, they radiate during the acceleration and deceleration processes around the spot position (approximately 0.5 $\mu\mathrm{m}$) with $\gamma$-photons emitted in the direction parallel to the target surface. Afterwards, the electrons mostly return towards the target surface. Periodically, another group of electrons are injected into the interference pattern and the process is repeated.
@@
Fig. LABEL:fig:trajectory-c shows the 2D Fourier transform of $B_{z}$ slice (at $z=0$) at a time of $90$ fs. The figure shows the HHG along the specular direction due to the mechanism of relativistic oscillating mirror (ROM) [80] (see also review article [9] and references cited therein). Furthermore, the figure shows the recently discovered XUV relativistic instability-modulated emission (RIME) propagating along the surface-parallel direction [16]. During the RIME process a strong quasi-static magnetic field at the target surface is generated, which confines part of the electrons near the plasma surface and prevents them from passing through it. In turn, oscillating relativistic nanobunches are generated via the two-stream instability, which are responsible for the surface-parallel XUV emission seen in Fig. LABEL:fig:trajectory-c, and which also help with injection of electrons from the target surface into the interference pattern in front of it. The presence of the quasi-static field is shown by the straight lines in Fig. LABEL:fig:trajectory-c in the area $k_{y}<0$. The straight lines disappear from the Fourier transform result when $B_{z}$ at the target position and beyond is artificially removed from the data, as is shown in the inset of Fig. LABEL:fig:trajectory-c.
@@
The $\gamma$-photon energy spectra at time of 60 fs (the time used in Figs. LABEL:fig:field_electron_gamma and LABEL:fig:trajectory) and at 100 fs (end of the simulation) for normal and ${}^{\circ}$incidence are compared in Fig. LABEL:fig:spectra_rate-a.

### v2

@@ Abstract
The interaction of a high-power laser with a solid target provides ways to produce beams of $\gamma$-photons. For normal incidence of the laser on the target the beams usually appear in a form of two lobes, which are symmetric with respect to the laser propagation axis. In this work we demonstrate via three-dimensional particle-in-cell simulations a regime where for oblique incidence the emission of a collimated $\gamma$-photon beam is in the direction parallel to the target surface. The process is ascribed to the interference pattern in the electromagnetic field formed by the incident and reflected laser pulse. The electromagnetic field accelerates electrons to the GeV energy level, while temporarily directing their momentum along the target surface. Consequently, they emit a collimated $\gamma$-photon beam in the same direction. The dependencies of $\gamma$-photon emission on the incident angle, laser pulse polarization, power and duration and target thickness and preplasma are also addressed in the paper. The beam directionality is important for designing future experiments. In addition, this setup causes the generation of high-order harmonics propagating along the target surface.
  • March 10, 2025
@@ 1 Introduction
An enormous progress in the development of extremely high-power laser technology resulting in the construction of many multi-PW laser facilities has been made all over the world [1]. As a result, there is a fast growth of interest towards building compact laser-based accelerators of charged particles [2, 3, 4, 5, 6, 7, 8] and radiation sources [9, 10]. Recent experiments on laser electron acceleration demonstrated electrons accelerated up to the energy of 10 gigaelectron volts (GeV) within a distance of tens of centimeters [11, 12]. In ion acceleration, proton beams with energy reaching 150 MeV have been demonstrated in the interaction of a laser with a thin foil target [13]. High-power lasers are used for studying high-brightness high-energy photon sources [14] and high-order harmonic generation (HHG) in the laser interaction with gas [15] and solid density targets [9, 16]. These lasers pave the way towards studying the previously unexplored domain of parameters corresponding to superstrong field for the quantum regimes of electromagnetic wave interaction with charged particles and the vacuum [17, 18, 19, 20]. These processes have been demonstrated through the electron-positron ($e^{-}$$e^{+}$) pair creation from vacuum in the seminal experiment [21]. A further experiment has been recently proposed for proving the nonlinear property of vacuum [22]. It was also predicted theoretically that through this interaction the $\gamma$-flash can be generated via multiphoton Compton scattering with high conversion efficiency [23, 24, 25, 26] and the first experimental evidence for this has been recently presented [27]. The multiphoton Compton scattering is characterized by the Lorentz invariant parameter $\chi_{e}=\gamma_{e}\sqrt{\left(\textbf{E}+\textbf{v}\times\textbf{B}\right)^{2}-\left({\textbf{v}\cdot\textbf{E}/c}\right)^{2}}$, where $\gamma_{e}$ is the Lorentz factor of the electron with velocity v, $c$ is the speed of light in vacuum, and E and B are the laser electric and magnetic fields normalized to the Schwinger field, $E_{S}\approx 1.3\times 10^{18}$ $\mathrm{V/m}$. The $\gamma$-flash radiation has exciting applications in various fields, such as materials science at extreme energy densities [28], $\gamma$-ray inspection and imagining [29], photonuclear reactions [30, 31, 32], neutron sources [33], photonuclear fission [34, 35], radiotherapy [36], shock-wave studies [37], $e^{-}$$e^{+}$ pair generation [38] (which can be further guided using orthogonal collision with a laser [39, 40]) and in quantum technologies[41], and it can be used to help understand the mechanisms of high-energy astrophysical processes [42, 43, 44, 45].
@@
A typical approach towards realization of $\gamma$-photon radiation uses the interaction of a laser pulse with a solid density target [46, 47, 48, 26]. Here, the laser pulse accelerates electrons, which then interact with the same pulse. Another setup to increase the $\gamma$-photon emission is to use two laser pulses [49, 50, 51, 52, 53] (or more of them [54, 55]). The normal incidence of a linearly polarized laser on target results in a symmetric double-lobe structure of the emitted $\gamma$-photons, which is demonstrated in particle-in-cell (PIC) simulations for various cases [24, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. However, in typical laser-solid interaction experiments, an oblique incidence angle is required to mitigate laser back-reflection which could damage the laser. In contrast to the double-lobe of the normal incidence, the structure and directionality of the emitted $\gamma$-photon beam(s) for oblique incidence varies strongly for different configurations. In addition to the incidence angle, the $\gamma$-photon directionality also depends on the target density, surface modulation [66] and target thickness, and various laser parameters (e.g., duration and focal spot size) [67]. Thus, the prediction of $\gamma$-photon beam direction becomes crucial for the reliable measurement of future experiments. A possible solution is to use specific laser-target parameters to align the dominant direction of the emitted $\gamma$-photons to the target surface. This was previously reported using grazing laser incidence in combination with a relativistically near-critical target [68] (although limiting the $\gamma$-photon yield) and in studies on lower-energy radiations like extreme ultraviolet (XUV) for oblique incidence (with a flat target [16] and a grating target [69]) and hard x-ray for parallel incidence [70]. The use of a grating on the target surface or a very grazing laser incidence angle are connected with generation of surface plasma waves which can be used for electron acceleration [71] and HHG [69, 70]. Moreover, for oblique incidence the interference between incident and reflected parts of the laser pulse generates a pattern in the electromagnetic field, accelerating electrons to high energies, as demonstrated via grazing incidence [72, 73]. This approach can be extended to a non-grazing incidence and opaque target (where the generated $\gamma$-flash is separated from the reflected laser pulse) as demonstrated in this work. The electron momentum is temporarily directed along the target surface by the field pattern, which results in a collimated $\gamma$-photon beam being emitted in the same direction.
@@
The paper is structured as follows. The theory of electron dynamics in the interference pattern of the electromagnetic field is described in Section 2. The simulation method and parameters are described in Section 3. The results are presented in Section 4, which is divided into two subsections. In subsection 4.1 we compare the normal incidence with ${}^{\circ}$oblique incidence and describe the mechanism behind the collimated $\gamma$-photon emission in the direction parallel to the target surface. This section also includes a Fourier transform analysis to observe the HHG. The dependence of the $\gamma$-photon emission on the laser incidence angle is shown in subsection 4.2. Section 5 contains the discussion of the results and the conclusion. The Appendix describes the dependency of $\gamma$-photon emission on laser polarization, pulse duration, laser power, target thickness and preplasma and the use of Virtual Reality to visualize our data.
@@ 3 Simulation setup and parameters
The 3D simulations were done with the PIC code EPOCH [74], including the Higuera-Cary pusher and QED module for non-linear Compton scattering. The Bremsstrahlung process [75, 76] is omitted in this work, as Compton scattering dominates over it at ultra-high laser intensities for times comparable to the laser pulse duration. The simulation box size is 14.336 $\mu$m $\times$ 20.48 $\mu$m $\times$ 10.24 $\mu$m, the cell size is 4 nm $\times$ 4 nm $\times$ 32 nm and 4 macro-electrons and 4 macro-ions are assigned to each cell containing plasma. A convergence test using double number of particles has been made with minimal changes observed. In all simulations the laser propagates from the boundaries at positions $x=y=-10.24$ $\mu$m at an angle of 45 (the plane of incidence is the $x$-$y$ plane), as shown in figure LABEL:fig:sim_setup. The incidence angle is determined by the target orientation, as shown in Fig. LABEL:fig:sim_setup(a) for the normal incidence and LABEL:fig:sim_setup(b) for the 45 incidence.
@@
The reference laser is linearly p-polarized. The laser wavelength is $0.8\ \mathrm{\mu m}$ and the peak intensity is $I_{\mathrm{max}}=2\times 10^{23}\ \mathrm{W/cm^{2}}$, thus $a_{0}\approx 0.85\sqrt{I\left[10^{18}\mathrm{W\ cm^{2}}\right]\lambda^{2}\left[\mathrm{\mu m}\right]}\approx 304$. The critical plasma density is $n_{c}=~{}m_{e}\omega^{2}/4\pi e^{2}\approx 1.742\times 10^{21}\ \mathrm{cm^{-3}}$, where $\omega$ is the laser angular frequency, $m_{e}$ is the electron mass and $e$ is the electron charge. The laser pulse has Gaussian spatial and temporal profiles with beam width equal to 3 $\lambda$ (2.4 $\mu$m) at the full width at half maximum (FWHM) and duration of 20 fs at FWHM (in intensity). The total laser energy for the reference case is 283 J in the 3D simulation, thus the laser power is approximately 14 PW. The target is made of fully ionized iron with an electron number density of $2.22\times 10^{24}\ \mathrm{cm}^{-3}$ (about 1274 $n_{c}$). The target thickness is 1.5 $\mu$m.
The QED module in EPOCH code includes the recoil effect on electrons where they lose momentum equal to the momentum of the emitted photons. Moreover, it provides options to include photons dynamics after they are generated and the production of $e^{-}$$e^{+}$ pairs. Aside of the possible pair production, the photons currently do not further interact with the plasma or electromagnetic fields and propagate ballistically after being generated. However, the effect of pair production is not significant for our simulation parameters as only one positron macroparticle (corresponding to 284160 positrons) is generated for the ${}^{\circ}$case (in comparison $2.5\times 10^{9}$ electron macroparticles of the same numerical weight is initialized in the simulation) and 15 positron macroparticles (corresponding to $4\times 10^{6}$ positrons) for the ${}^{\circ}$case of simulations shown in Fig. LABEL:fig:sim_setup. Therefore, we disable photons dynamics and the pair production afterward to reduce the computational time. Furthermore, only $\gamma$-photons and electrons with energy above 1 MeV are stored for evaluation (photons with lower energy are still generated and contribute to the electron recoil). The visualization of the $\gamma$-photon emission temporal evolution of the ${}^{\circ}$case can be accessed online [77] via our Virtual Beamline – VBL application [78, 79, 80], described at the end of the Appendix.
@@ 4.1 Generation of collimated $\gamma$-photon beam along the target surface
The $\gamma$-photon directionality is shown in Fig. LABEL:fig:spherical_distribution_00x45 via the angular distribution of the emitted $\gamma$-photon energy in the 3D space. Due to the setup geometry with the $x$-$y$ polarization plane (whose 2D slices at $z=0$ are further examined later), it is convenient to represent the angular distribution using the angles $\theta=\mathrm{arctan}(p_{y}/p_{x})$ and $\phi=\mathrm{arctan}(p_{z}/p_{r})$ in Fig. LABEL:fig:spherical_distribution_00x45, where $p_{x,y,z}$ are the momenta in the respective directions of the Cartesian system of the simulation and $p_{r}=\sqrt{p_{x}^{2}+p_{y}^{2}}$. Note that as targets are tilted by a different angle for different laser incidence angles, the angular data are shifted in the $\theta$-direction by an angle equal to the angle between the target and the $y$-axis of the simulation. The normal incidence case (Fig. LABEL:fig:spherical_distribution_00x45-a) creates a typical double-lobe structure. For our simulation parameters the maximum energy density is observed around ±$\pm 50^{\circ}$. The situation differs for the 45 case (figure LABEL:fig:spherical_distribution_00x45-b) where a single spot around 90 (direction parallel to the target surface) is observed. The lineouts of the blue lines (right part of Fig. LABEL:fig:spherical_distribution_00x45) show the energy distribution in the transverse direction at angles $50^{\circ}$ and $90^{\circ}$ for the normal and ${}^{\circ}$incidence cases respectively. In both cases the $\gamma$-photon beam is well collimated in the transverse direction (angular spread at FWHM is around ${}^{\circ}$for both cases) but the maximal energy density is more than double in the case of oblique incidence. The brilliance in Fig. LABEL:fig:spherical_distribution_00x45-c is calculated at the direction of the peak $\gamma$-photon yield (at the same angles as the lineouts) using an opening full-angle of 10. Our simulations reveal a high brilliance for the ${}^{\circ}$case, similar to that of Ref. [63], approaching $10^{23}$ s-1mm-2mrad-2 per 0.1% bandwidth in the tens of MeV $\gamma$-photons energy range. In comparison, the brilliance for the normal incidence case is approximately three times lower, which is reflected to the ratio of the conversion efficiency, as discussed in Section 4.2.
@@
For the $45^{\circ}$ laser incidence case (shown in Fig. LABEL:fig:field_electron_gamma-b,d) $\gamma$-photons are emitted by three distinct mechanisms, occurring at different locations and with different electron trajectory patterns.
Firstly, as described in Section 2, the reflected laser pulse interferes with the incident one and creates an interference pattern, which propagates with superluminal phase velocity along the target surface (as seen in Fig. LABEL:fig:field_electron_gamma-b). Inside this pattern, electrons are accelerated to energies several times higher than that given by the quiver velocity [73]. As the field pattern moves faster than electrons, they experience a strong acceleration (and subsequent deceleration when the pattern changes sign) while temporarily aligning their momentum with the target surface. The direction of motion of the emitted $\gamma$-photons is approximately the same as that of the emitting electrons. Therefore, directional $\gamma$-photon emission occurs at the field extremes, as in Figs. LABEL:fig:field_electron_gamma-b,d. The field extremes and the regions of high $\gamma$-photon energy density appear at $y=[-1.7,-0.6,0.55,1.7]$ $\mu\mathrm{m}$, with their spacing as predicted by the equation $\lambda_{\perp}=\lambda/\sin\theta_{n}=0.8\ \mu\mathrm{m}/\sin(45^{\circ})\approx 1.13\ \mu\mathrm{m}$. These spots appears around $x=0$, which at time 60 fs is a low-density plasma region in front of the target (as the target surface shifts inwards under the radiation pressure). The instantaneous target surface is shown by the green contour in Fig. LABEL:fig:field_electron_gamma.
@@
Thirdly, with an even smaller contribution, a $\gamma$-photon population is emitted from the instantaneous target surface ($y=[-1,0.25]$ $\mu\mathrm{m}$ at 60 fs). There, electrons propagate along the target surface with wiggling movement, as seen in Fig. LABEL:fig:field_electron_gamma-b. Note that in our case the electron population propagating along the target surface (as in the surface plasma wave [69, 71, 70]), moving outside of the interference pattern region ($y\approx\pm 3\mu$m) does not significantly contribute to the $\gamma$-photon emission as the $\chi_{e}$ factor for those electrons is small.
@@
Here, $C=e^{2}m_{e}^{2}/3\sqrt{3}\pi\hbar^{2}$, where $\hbar$ is normalized Planck constant, $K_{2/3}$ is the Modified Bessel function of the second kind and $u$ is the integration variable. Note that Fig. LABEL:fig:trajectory-a is similar to Fig. LABEL:fig:field_electron_gamma-d (although Fig. LABEL:fig:trajectory-a is more clear since the electron data are instantaneous).
@@
Next, we tracked electrons with $\chi_{e}\geq 0.3$ (34 macroparticles), located in the bright spot area shown in Fig. LABEL:fig:trajectory-a. The electron trajectories are shown in Fig. LABEL:fig:trajectory-b, where their energy is represented by the color scale (time evolution of the location of selected particles in $B_{z}$, $E_{x}$, $E_{y}$ is shown in the supplementary video). The evolving (with respect to the $y$-position) front target surface is depicted by the green line, the $x$-position represents the first position with electron number density equal or higher than the initial target density at the corresponding $y$-position and $z=0$, where the $y$-positions are linked to one of the tracked electrons. The tracked electrons begin with a wiggling movement close to the front target surface as they reflect from the strong quasi-static magnetic field (generated by the space charge inhomogeneity). Afterwards, they are injected into the field interference pattern, align parallel to the initial target surface, and are accelerated to energies up to 700 MeV. Consequently, they radiate during the acceleration and deceleration processes around the spot position (approximately $y=-0.5$ $\mu\mathrm{m}$) with $\gamma$-photons emitted in the direction parallel to the target surface. Afterwards, the electrons mostly return towards the target surface. Periodically, another group of electrons are injected into the superluminal interference pattern and the process is repeated.
@@
Fig. LABEL:fig:trajectory-c shows the 2D Fourier transform of $B_{z}$ slice (at $z=0$) at a time of $90$ fs. The figure shows the HHG along the specular direction due to the mechanism of relativistic oscillating mirror (ROM) [82] (see also review article [9] and references cited therein). Furthermore, the figure shows the recently discovered XUV relativistic instability-modulated emission (RIME) propagating along the surface-parallel direction [16]. The RIME process is driven by a strong quasi-static magnetic field at the target surface which is generated by the oblique incidence [83]. This field confines part of the electrons near the plasma surface and prevents them from passing through it. In turn, oscillating relativistic nanobunches are generated via the two-stream instability, which are responsible for the surface-parallel XUV emission seen in Fig. LABEL:fig:trajectory-c, and which also help with injection of electrons from the target surface into the interference pattern in front of it. The presence of the quasi-static field is shown by the straight lines in Fig. LABEL:fig:trajectory-c in the area $k_{y}<0$. The straight lines disappear from the Fourier transform result when $B_{z}$ at the target position and beyond is artificially removed from the data, as is shown in the inset of Fig. LABEL:fig:trajectory-c.
@@
The $\gamma$-photon energy spectra at time of 60 fs (the time used in Figs. LABEL:fig:field_electron_gamma and LABEL:fig:trajectory(a)) and at 100 fs (end of the simulation) for normal and ${}^{\circ}$incidence are compared in Fig. LABEL:fig:spectra_rate-a.